OpenStudy (shamil98):

A question about implicit differentiation.. x^3 + y^3 - 9xy = 0

OpenStudy (shamil98):

\[\large \frac{ d }{ dx } x^3 + \frac{ d }{ dx }y^3 - \frac{ d }{ dx } 9xy = \frac{ d }{ dx }0\] I know how to derive the x^3 and y^3 and 0 ofc.. but I don't understand what I am suppose to with the -9xy..

OpenStudy (anonymous):

use product rule d(xy)= x.dy+y.dx

OpenStudy (shamil98):

Oh. That makes sense.

OpenStudy (shamil98):

so d/dx (-9xy) = -9y + -9x(1)?..

OpenStudy (anonymous):

no, d/dx(-9xy) = -9x (dy/dx) - 9y

OpenStudy (anonymous):

actually is d(x)/dx (-9y) + d(y)/dx(-9x)

OpenStudy (anonymous):

then become 1(-9y)+dy/dx(-9x)

OpenStudy (anonymous):

then become -9y-dy/dx(9x)

OpenStudy (shamil98):

Why is the dy/dx still there though? ..

OpenStudy (anonymous):

because chain rule state tht d/dx(xy) = dx/dx (y)+dy/dx(y)

OpenStudy (anonymous):

sry is product rule... XD

OpenStudy (anonymous):

so when ever u hv 2 variable, u will need to apply this product rule

OpenStudy (shamil98):

it's fine lol. i'm still a bit lost though. According to my book it becomes.. \[\large 3x^2 + 3y^2 \frac{ d }{ dx } - 9 (x \frac{ dy }{ dx } + y \frac{ dx }{ dx }) = 0\] Couldn't x dy/dx be written as xy' ?

OpenStudy (anonymous):

x (dy/dx) and xy' is the same things, but 2nd term in not correct, when u hv d/dx(3y^3) u need to take care, i will become like this, dy/dx*d/dy(3y^3)

OpenStudy (anonymous):

sry is dy/dx*d/dy(y^3), then (dy/dx)(3y^2)

OpenStudy (anonymous):

remember this, \[\frac{ d }{ dx }=\frac{ dy }{ dx }\times \frac{ d }{ dy }\]

OpenStudy (anonymous):

see is a chain rule, when u cancel out the dy u will get back d/dx

OpenStudy (shamil98):

leibniz notation always confuses me lol.

OpenStudy (anonymous):

erm, i get use to this notation... ><, do u clear abt this?? or still blur?

OpenStudy (anonymous):

when u r dealing with implicit differentiation, u go to term by term, so tht u will be easier, for checking and solving.

OpenStudy (anonymous):

x^3 + y^3 - 9xy = 0 first term d/dx(x^3)=3x^2 second term d/dx(y^3)=(dy/dx)(d/dy)(y^3)=(dy/dx)3y^2 third term d/dx(-9xy)=-9(d/dx(xy))=-9(x(dy/dx)+y(dx/dx))=-9(x(dy/dx)+y) overall, 3x^2 + (dy/dx)3y^2 - -9(x(dy/dx)+y) = 0

OpenStudy (shamil98):

I think I got it . \[\large \frac{ d }{ dx } (x^3+y^3-9xy = 0)\] \[\large 3x^2 + 3y^2y' - 9xy' - 9y = 0\] Group the y' together and get terms on one side. \[\large y'(3y^2 - 9x) = -3x^2 + 9x\] Divide. \[\large y' = \frac{ 3(-x^2+3x) }{ 3(y^2-3x) } = \frac{ -x^2+3x }{ y^2-3x }\]

OpenStudy (anonymous):

yup, urs correct, my equation hv some error on 3rd term is 1 minus sign only...

OpenStudy (shamil98):

Oops, I made a mistake i accidentally made the -9y into 9x instead of 9y*

OpenStudy (shamil98):

\[\large y' = \frac{ 3y - x^2 }{ y^2 - 3x }\] there we go.

OpenStudy (anonymous):

yup =)

OpenStudy (shamil98):

Thanks.